### v2

# Robust Control Lyapunov-Value Functions for Nonlinear Disturbed Systems

Abstract
This article presents a method to construct robust control Lyapunov value functions (R-CLVFs) for robust and stabilizing control of nonlinear systems with input and disturbance bounds. Through modifications to Hamilton-Jacobi reachability analysis, the R-CLVF can be computed via dynamic programming. We prove that the R-CLVF can be used to stabilize the system to its smallest robust control invariant set at a user-specified exponential rate. The R-CLVF additionally can be used to obtain the domain over which stabilization at the desired rate is possible, i.e., the region of exponential stabilizability. Due to the computationally expensive nature of dynamic programming, we additionally propose methods to improve tractability via decomposing the dynamical model or warm-starting the computation. Under certain assumptions, we show that these approaches maintain exact solutions. Three numerical examples are provided, validating our definition of the smallest robust control invariant set, illustrating the impact of the exponential rate and different loss functions, and demonstrating the efficiency of computation using warm-starting and decomposition.
I INTRODUCTION
Liveness and safety are two main concerns for autonomous systems working in the real world. Using control Lyapunov functions (CLFs) to stabilize the trajectories of a system to an equilibrium point [1, 2, 3] is a popular approach to ensure liveness, whereas using control barrier functions (CBFs) to guarantee forward control invariance is popular for maintaining safety [4, 5, 6]. However, identifying valid CLFs and CBFs is challenging for general nonlinear systems, and users of these methods typically rely on hand-designed or application-specific CLFs and CBFs [7, 8, 9, 10, 11]. Finding these hand-crafted functions can be difficult, especially for high-dimensional systems with state and/or input constraints.
Liveness and safety can also be achieved by formal methods such as Hamilton-Jacobi (HJ) reachability analysis [12]. This method formulates liveness and safety as optimal control problems, and has been used for applications in aerospace, autonomous driving, and more [13, 14, 15, 16, 17]. This method computes a value function whose level sets provide information about safety (or liveness) over space and time, and whose gradients inform the safety (or liveness) controller. This value function is the unique viscosity solution to a Hamilton-Jacobi-Issac’s Variational Inequality (HJI-VI), can be computed numerically using dynamic programming (DP) for general nonlinear systems, and can accommodate input and disturbance bounds. Undermining these appealing benefits is the curse of dimensionality: computation scales exponentially with state dimension. Ongoing research has improved computational efficiency [18, 19, 20, 21], but performing DP in high dimensions (6D or more) remains challenging.
Standard HJ reachability analysis focuses on problems such as the minimum time to reach a goal, or how to avoid certain states for a finite (or infinite) time horizon. It does not stabilize a system to a goal after reaching it. In our previous work [22], we modified the value function and defined the control Lyapunov value function (CLVF) for undisturbed nonlinear systems with bounded control. The CLVF finds the smallest control invariant set (SCIS) and the region of exponential stabilizability (ROES) of the system given a user-specified exponential rate $\gamma$. Its gradient can be used to synthesize controllers that stabilize the system to the SCIS at the rate $\gamma$. The stabilizability comes from the fact that the CLVF is the viscosity solution to a modified variational inequality. However, there is one technical error in the previous work: the CLVF is not the unique viscosity solution to the VI, and we provide a counterexample in this article. The original work [22] is restricted in several aspects: it considers only systems without disturbance, the choice of the loss function is restricted to 2-norm, the curse of dimensionality restricts its application to relatively low-dimensional systems (5D or lower), and some concepts would benefit from deeper discussion.
Other works that study the relationship between CLF and some partial differential equations (PDE) include Zubov’s method and its extensions [23, 8, 24, 25]. Since the solution to the PDE is generally not continuously differentiable, the differential inclusion is usually considered. Another work that considers the relation between the state-constrained optimal control problem and the HJI-VI is [26], where an auxiliary unconstrained optimal control problem is solved using the HJI-VI. However, this requires augmenting to one additional dimension, which requires much larger computational resources.
This work seeks to address all the restrictions of the previous CLVF work, and provide more rigorous proofs of the theorems. The main contributions are:
  1. 1.
    We define the time-varying robust CLVF (TV-R-CLVF) and the robust CLVF (R-CLVF) for systems with bounded disturbance and control. We prove that the R-CLVF is Lipschitz continuous, satisfies the dynamic programming principle (DPP), and is the viscosity solution to the corresponding R-CLVF variational inequality (VI). The algorithm for computing the R-CLVF is updated.
  2. 2.
    We define the smallest robustly control invariant set (SRCIS) of a point of interest (POI) given system dynamics and prove that the SRCIS is the zero sub-level set of the R-CLVF. We show that the R-CLVF can stabilize the system to the SRCIS of the POI from the ROES with an exponential rate $\gamma$ defined by the user.
  3. 3.
    Two methods to accelerate computation are introduced: warm-starting and system decomposition. We prove that under certain assumptions, these acceleration methods will recover exact results.
  4. 4.
    A point-wise optimal R-CLVF quadratic program (QP) control law is provided, which is guaranteed to be feasible, and numerical examples are provided to validate the theory and show numerical efficiency with warm-starting and system decomposition.
The paper is organized in the following order: Section II provides background information on HJ reachability analysis and CLVF. Section III introduces the TV-R-CLVF, and builds up the theoretical foundation for the R-CLVF. A feasibility-guaranteed quadratic program controller is provided. Section IV introduces warm-starting and system decomposition to accelerate the computation. Section V shows three numerical examples, validating the theory.
II BACKGROUND
In this paper, we seek to exponentially stabilize a given nonlinear time-invariant dynamic system with bounded control and disturbance to its SRCIS. We start by defining crucial some terms.
II-A Problem Formulation
Consider the nonlinear time-invariant system
$\dot{x}(s)=f\left(x(s),u(s),d(s)\right),\hskip 5.0pts\in[t,0],\hskip 5.0ptx(t)=x_{0},$ (1)
where $t<0$ is the initial time, and $x_{0}\in\mathbb{R}^{n}$ is the initial state. The control signal $u(\cdot)$ and disturbance signal $d(\cdot)$ are drawn from the set of measurable functions $\mathbb{U}$ and $\mathbb{D}$. Assume also the control input $u$ and disturbance $d$ are drawn from convex compact sets $\mathcal{U}\subset\mathbb{R}^{m}$ and $\mathcal{D}\subset\mathbb{R}^{p}$ respectively. We have:
$\mathbb{U}:=\{u(\cdot):[t,0]\mapsto\mathcal{U},u(\cdot)\text{ is measurable}\},$
$\mathbb{D}:=\{d(\cdot):[t,0]\mapsto\mathcal{D},d(\cdot)\text{ is measurable}\}.$
We make the following assumptions about the system:
  1. ($A_{1}$)
    The dynamic model $f:\mathbb{R}^{n}\times\mathcal{U}\times\mathcal{D}\mapsto\mathbb{R}^{n}$ is uniformly continuous in $(x,u,d)$, Lipschitz continuous in $x$ for fixed $u(\cdot)$ and $d(\cdot)$, with Lipschitz constant $L_{f}$.
  2. ($A_{2}$)
    The dynamic model $f:\mathbb{R}^{n}\times\mathcal{U}\times\mathcal{D}\mapsto\mathbb{R}^{n}$ is bounded $\forall x\in\mathbb{R}^{n},u\in\mathcal{U},d\in\mathcal{D}$.
Under these assumptions, given an initial state $x$ and control and disturbance signals $u(\cdot)$, $d(\cdot)$, there exists a unique solution $\xi(s;t,x,u(\cdot),d(\cdot))$, $s\in[t,0]$ of the system (1). We call this solution the trajectory of the system, and where possible, we use $\xi(s)$ for conciseness. Further assume the disturbance signal can be determined as a strategy with respect to the control signal: $\lambda:\mathbb{U}\mapsto\mathbb{D}$, drawn from the set of non-anticipative maps $\lambda\in\Lambda$  [27], defined as:
[…]
$\implies\mathcal{N}[a](s)=\mathcal{N}[\hat{a}](s)\text{ a.e. }\forall s\in[t,0]\}.$
In this paper, we seek to stabilize the system (1) to its SRCIS. We first introduce the notion of a robust control invariant set (RCIS).
Definition 1
(RCIS) A closed set $\mathcal{I}$ is robustly control invariant for (1) if $\forall x\in\mathcal{I}$, $\forall\lambda\in\Lambda$, $\exists u(\cdot)\in\mathbb{U}$ such that $\xi(s;t,x,u(\cdot),\lambda[u])\in\mathcal{I}$, $\forall s\in[t,0]$.
We also assume the following:
  1. ($A_{3}$)
    When the system has equilibrium points, the origin 0 is one, i.e. $f($0,0,0)=)=) = 0.
  2. ($A_{4}$)
    Given any point of interest (including the origin), there exists a robustly control invariant set, whose convex hull contains the point of interest.
Assumption $A_{3}$ is standard. For $A_{4}$, some systems might not posses an equilibrium point, e.g., the 3D Dubins car with constant velocity or any systems that do not satisfy the small control property.
Remark 1
$A_{4}$ is a weaker assumption compared to assuming the existence of a CLF. For a valid CLF, the region of null controllability (RONC) is also robust control invariant. Further, all trajectories starting from the RONC can be stabilized to the equilibrium point. This is considered as one key contribution of the paper: we extend the classical definition of stabilizing to the equilibrium point to stabilizing to the SRCIS, if the former is impossible to achieve.
We are also interested in finding the ROES of a set. We first define the signed distance from a point to a set $\mathcal{A}$ to be
$\text{dst}(x;\mathcal{A})=\begin{cases}\min_{a\in\partial\mathcal{A}}||x-a||&x\notin\mathcal{A},\\ -\min_{a\in\partial\mathcal{A}}||x-a||&x\in\text{int}(\mathcal{A}),\\ 0&x\in\partial\mathcal{A}.\end{cases}$ (2)
where $\partial\mathcal{A}$ is the boundary of $\mathcal{A}$ and any vector norm is applicable here.
Definition 2
The ROES of a set $\mathcal{I}$ is the set of states from which the trajectory converges to $\mathcal{I}$ with an exponential rate $\gamma$:
[…]
$\text{dst}(\xi(s;t,x,u(\cdot),\lambda[u]);\mathcal{I})\leq ce^{-\gamma(s-t)}\text{dst}(x;\mathcal{I})\}.$
II-B HJ Reachability and SRCIS
In the previous work [22], we proposed constructing the CLVF using HJ reachability analysis. This is done by formulating a reachability safety problem, where the system tries to avoid all regions of the state space that are not the origin. This problem can be solved as an optimal control problem.
Traditionally in HJ reachability analysis, a continuous loss function $h:\mathbb{R}^{n}\mapsto\mathbb{R}$ is defined such that its zero super-level set is the failure set =$\mathcal{F}=\{x:h(x)>0\}$. The finite-time horizon cost function captures whether a trajectory enters $\mathcal{F}$ at any time in $[t,0]$ under given control and disturbance signals by computing the maximum loss accrued over time:
$J(t,x,u(\cdot),d(\cdot))=\max_{s\in[t,0]}h\bigl{(}\xi(s;t,x,u(\cdot),d(\cdot))\bigl{)}.$ (3)
The value function is the cost under the optimal control signal and worst-case disturbance:
$V(x,t)$ $=\sup_{\lambda\in\Lambda}\inf_{u(\cdot)\in\mathbb{U}}J(t,x,u(\cdot),\lambda[u])$
$=\sup_{\lambda\in\Lambda}\inf_{u(\cdot)\in\mathbb{U}}\max_{s\in[t,0]}h(\xi(s;t,x,u(\cdot),\lambda[u]).$ (4)
For a given $t$, the (strict) zero super-level set $\mathcal{V}_{0}=\{x:V(x,t)>0\}$ denotes the set of initial states such that there exists a disturbance signal that drives the trajectory to $\mathcal{F}$ for some time $s\in[t,0]$, despite the control signal used. The zero sub-level set of $V(x,t)$ is therefore safe for the time horizon $[t,0]$. This can be extended to say that each $\alpha$ sub-level set $\mathcal{V}_{\alpha}=\{x:V(x,t)\leq\alpha\}$ is safe w.r.t. the set defined by $\mathcal{F}_{\alpha}=\{x:h(x)>\alpha\}$.
The infinite-time horizon value function is defined by taking the limit (if it exists) of $V(x,t)$ as $t\rightarrow-\infty$ [28],
$V^{\infty}(x)=\lim_{t\rightarrow-\infty}V(x,t).$ (5)
Different from the time-varying value function (II-B), for all states in the $\alpha$ sub-level set of $V^{\infty}(x)$, there always exists a control signal such that the maximum loss is lower than $\alpha$ despite the disturbance signal. This means every $\alpha$ sub-level set of $V^{\infty}(x)$ is robustly control invariant, and the trajectories can be maintained within a particular level set boundary. Further, this set is the largest RCIS contained within the $\alpha$ sub-level set of $h(x)$ [29].
It has been shown that (II-B) is the unique viscosity solution to the following HJI-VI [30]:
$0=\min\biggl{\{}h(x)-V(x,t),$ (6)
$\hskip 10.00002ptD_{t}V(x,t)+\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}V(x,t)\cdot f(x,u,d)\biggl{\}},$
where $D_{t}$ and $D_{x}$ denote the derivative over time and state respectively. The value function (II-B) can be computed numerically using DP by solving this HJI-VI recursively over time. And the infinite-time value function (5) can be obtained by solving this HJI-VI until convergence.
Remark 2
In this paper, we restrict the selection of $h(x)$ to be vector norms (e.g., p-norms, or weighted Q norms,) and the same as the norm used in equation (2). More specifically, given any POI $p\in\mathbb{R}^{n}$, we pick $h(x;p)=\|x-p\|$. In other words, the loss function measures the distance of a state to the POI. With this restriction, the cost function (3) captures the largest deviation from the POI of a given trajectory, initialized at $x$ with $u(\cdot)$ and $d(\cdot)$ applied, in time horizon $[t,0]$. The infinite time value function (5) captures the largest deviation with optimal control and disturbance signals applied in an infinite time horizon.
Denote the minimal value of $V^{\infty}(x)$ as $V^{\infty}_{m}:=\min_{x}V^{\infty}(x)$. The $V^{\infty}_{m}$-level set of $V^{\infty}$ is the smallest RCIS (SRCIS), and denoted by $\mathcal{I}_{m}$. Different norms result in different SRCISs.
Definition 3
(SRCIS.) A closed set $\mathcal{I}_{m}(p)$ is the SRCIS of a given POI $p$ and norm chosen if it is the level set corresponding to the minimal value of $V^{\infty}$.
Since the SRCIS is the level set corresponding to the minimal value, all states in it have the same value $V^{\infty}_{m}$, and we have
$V^{\infty}_{m}=\max_{x\in\mathcal{I}_{m}}h(x;p).$ (7)
In other words, any trajectory starting from $\mathcal{I}_{m}$ has the same largest deviation along time (measured by $h$) to the POI, which is the value $V^{\infty}_{m}$.
Remark 3
The SRCIS should be understood as ‘the RCIS, with the smallest largest deviation to the origin.’ Here the term smallest means the ‘smallest level’ of $V^{\infty}$, and largest means the largest deviation along time. Note that there may be other RCISs contained in the SRCIS, but they are not level sets of $V^{\infty}$. This is different from the ‘minimal RCIS’ as defined in [31, 32], where ‘minimal’ is defined as ‘no subset is robust control invariant’. For illustration, consider the following example:
$\dot{x}_{1}=-x_{1}+d,\quad\dot{x}_{2}=x_{2}+u$ (8)
where $u\in[-1,1]$ and $d\in[-0.5,0.5]$. This system has an undisturbed, uncontrolled equilibrium point $[x,u,d]=[\textbf{0},0,0]$. It can be verified that =$\mathcal{I}=\{x_{1}\in[-0.5,0.5],x_{2}=0\}$ is one ‘minimal RCIS’ as all its subsets are not robustly control invariant. In fact, picking any $x_{2}\in[-1,1]$ results in a ‘minimal RCIS.’ On the other hand, picking $h(x)=||x||_{\infty}$, the SRCIS is $\mathcal{I}_{m}=\{x_{1},x_{2}\in[-0.5,0.5]\}$. This is because though the control can stabilize any $|x_{2}|<1$ to the origin, the disturbance is also strong enough to perturb any $|x_{1}|<0.5$ to leave the origin. Therefore, all states s.t. $x_{1},x_{2}\in[-0.5,0.5]$ have the same value, and the SRCIS measured by the $\infty$-norm is a square. Fig. 1 shows the SRCIS for three different choices of $h(x)$ and the corresponding value function.
Figure 1: SRCIS corresponds to different loss functions for system (8). Top left to right: different loss functions, including the 2-norm, infinity norm, and weighted Q-norm ($Q=diag[0.2,1]$). Middle left to right: R-CLVF ($\gamma=0$) when $h(x)=||x||_{2}$, $||x||_{\infty}$, $||x||_{Q}$, and $||x||_{Q}=\sqrt{x^{T}Qx}$. Bottom left to right: the corresponding SRCIS and a trajectory starting inside the SRCIS. The robust control invariance is validated.
An interesting observation is that adding or subtracting a constant value to the loss function $h(x)$, the corresponding SRCIS stays the same.
Proposition 1
Define $\underline{h}(x)=h(x)-a$, and denote the corresponding value function as $\underline{V}(x,t)$, we have $\underline{V}(x,t)=V(x,t)-a$.
Proof:
$\underline{V}(x,t)$ $=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max_{s\in[t,0]}\underline{h}(\xi(s;t,x,u(\cdot),\lambda[u])$
$=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max_{s\in[t,0]}\big{(}h(\xi(s;t,x,u(\cdot),\lambda[u])-a\big{)}$
$=V(x,t)-a,$ (9)
where the last step is because $a$ is a constant. ∎
Also note that we dropped $p$ from $h$ in Prop. 1, because it does not affect the result.
Each level set of the HJ value function (5) is only robustly control invariant, there is no guarantee that the system can be stabilized to lower level sets or the origin. In our preliminary work [22], we define the CLVF for undisturbed systems. In this article, we further develop the theory of CLVFs for disturbed systems and try to stabilize the system to any arbitrary POI. We also provide the necessary theorems for numerical implementation in high-dimensional nonlinear systems.
III ROBUST CONTROL LYAPUNOV-VALUE FUNCTIONS
In this section, we start by defining the time-varying robust CLVF (TV-R-CLVF) and prove some important properties of it. We then define the R-CLVF, which is the limit function of the TV-R-CLVF. We show that the existence of the R-CLVF is equivalent to the exponential stabilizability of the system to its SRCIS of a POI and that its domain is the ROES. This section is organized as follows: Section III-A defines the TV-R-CLVF and shows some key properties of it. Section III-B defines the R-CLVF and shows it satisfies the DPP (Theorem 3) and is the viscosity solution of a variational inequality (Theorem 4). Section III-C shows the sufficient and necessary conditions for the existence of the R-CLVF (Theorem 7). Section III-D studies the impact of different $\gamma$. It shows for all non-negative $\gamma$, the SRCISs are all the same (Lemma 8). Further, a larger $\gamma$ corresponds to a faster convergence rate, while a smaller ROES. Section III-E provides an feasibility guaranteed QP controller.
III-A TV-R-CLVF
Definition 4
A TV-R-CLVF is a function $V_{\gamma}(x,t):\mathbb{R}^{n}\times\mathbb{R}_{-}\rightarrow\mathbb{R}$ defined as:
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u(\cdot)\in\mathbb{U}}J_{\gamma}(t,x,u(\cdot),\lambda),$ (10)
where $J_{\gamma}(t,x,u(\cdot),d(\cdot))$ is the cost function:
$J_{\gamma}(t,x,u(\cdot),d(\cdot)))=\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\bigl{(}\xi(s;t,x,u(\cdot),d(\cdot))\bigl{)},$ (11)
$\gamma\geq 0$ is a user-specified parameter that represents the desired decay rate, and $\ell(x;p)=h(x;p)-V^{\infty}_{m}$.
The cost at a state captures the maximum exponentially amplified distance between the trajectory (starting from this state) and the zero-level set of $\ell(x)$. The optimal control tries to minimize this cost and seeks to drive the system towards the POI. In contrast, the disturbance tries to maximize the cost and push the system away from the POI.
From (10) and (11), if a trajectory is initialized outside the zero sub-level set of $\ell(x)$, the value $V_{\gamma}(x,t)$ will always be positive. If starting inside the zero sub-level set of $\ell(x)$, the value $V_{\gamma}(x,t)$ might be positive or negative, depending on the time horizon $t$. In other words, as $t$ decreases, the zero sub-level set of $V_{\gamma}(x,t)$ will shrink. In fact, we will show later that, given $t$, the zero sub-level set of $V_{\gamma}(x,t)$ is exactly the zero sub-level set of $V(x,t)-V^{\infty}_{m}$ (Lemma 8).
It should be noted that $\gamma$ serves as an exponential amplifier for the TV-R-CLVF, which in turn sets the desired exponential convergence rate of a trajectory to the SRCIS. This is introduced more formally in Lemma 5.
We now prove some properties of the TV-R-CLVF in Proposition 2 and show the mathematical foundation for how it can be obtained in Theorem 1 and 2.
Proposition 2
The TV-R-CLVF is Lipschitz in $x,t$ and bounded for any open set $\mathcal{C}$
Proof:
See the Appendix of [33]. ∎
We now present that the TV-R-CLVF satisfies the dynamic programming principle, and is the unique viscosity solution to the TV-R-CLVF VI.
Theorem 1
$V_{\gamma}(x,t)$ satisfies the following dynamic programming principle for all $t<t+\delta\leq 0$:
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}V_{\gamma}(\xi(t+\delta),t+\delta),$
$\hskip 60.00009pt\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}.$ (12)
Theorem 2
The TV-R-CLVF is the unique viscosity solution to the following VI,
$\max\biggl{\{}\ell(x)-V_{\gamma}(x,t),$ (13)
$\hskip 2.84544ptD_{t}V_{\gamma}+\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}V_{\gamma}\cdot f(x,u,d)+\gamma V_{\gamma}\biggl{\}}=0,$
with terminal condition $V_{\gamma}(x,0)=\ell(x)$.
The proofs of the above two Theorems can be obtained analogously following Theorems 2 and 3 in [29], and are omitted here. It should be noted that the definitions of viscosity solutions for the terminal value problem (TVP) and initial value problem (IVP) are slightly different. For more details, the reader is referred to [15, 16] for IPV and [34] for TVP.
With Theorems 1 and 2, the TV-R-CLVF can be again obtained by solving equation (13) for $s\in[t,0]$. Note that in equation (13), define $H:\mathcal{D}_{\gamma}\times\mathbb{R}^{n}\mapsto\mathbb{R}$,
$H(x,p)=\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}p\cdot f(x,u,d).$
$H$ is called the Hamiltonian, and the compactness of $\mathcal{U}$ and $\mathcal{D}$ and continuity of $f$ guarantee that $H$ is continuous w.r.t $(x,p)$.
III-B R-CLVF
In this part, we focus on the infinite-time horizon and introduce the R-CLVF.
Definition 5
(R-CLVF) Given an open set $D_{\gamma}\subseteq\mathbb{R}^{n}$ and $\gamma>0$, the function $V_{\gamma}^{\infty}:\mathcal{D}_{\gamma}\mapsto\mathbb{R}$ is a R-CLVF if the following limit exists:
$V_{\gamma}^{\infty}(x)=\lim_{t\rightarrow-\infty}V_{\gamma}(x,t).$ (14)
Notice the domain of the R-CLVF is $\mathcal{D}_{\gamma}$, where $\mathcal{D}_{\gamma}$ can be $\mathbb{R}^{n}$ if the limit exists on $\mathbb{R}^{n}$, while for the TV-R-CLVF, it is always $\mathbb{R}^{n}$. Also notice that for the R-CLVF, $\gamma$ is strictly greater than 0. The reason is that if $\gamma=0$, the limit in (14) may exists on a closed set.
Remark 4
Note that though in Definition 4 from [22], we claimed the domain is a compact set; unfortunately, this is incorrect. As a result, the proof in Remark 1 in [22] is also incorrect, and we can show that the convergence is not uniform. To see these, consider a 1D toy example $\dot{x}=x+u$, where $u\in[-1,1]$ and $\gamma>0$. Then, for any trajectory starting from $x\in(-1,1)$, there exists a control signal so that the trajectory reach 00 at some finite time, therefore the limit in (14) exists. For trajectories starting from $x=\pm 1$, with best control efforts, they can only stay at the same point, so the limit in (14) does not exist. For trajectories starting from $x>1$ or $x<-1$, the trajectory will diverge, and the limit in (14) does not exist. By taking feedback
$u^{*}(x)=1\text{ if }x<0,-1\text{ if }x>0,0\text{ if }x=0,$
$V_{\gamma}^{\infty}(x)$ is attained at $t=\min\{0,-\ln\frac{\gamma}{(1+\gamma)(1-|x|)}\}$, and as $x\rightarrow\pm 1$, $t\rightarrow-\infty$, so the convergence is not uniform. Further, we could compute the maximum value is actually given by $|x|$ if $t=0$ and $\frac{\gamma^{\gamma}}{(1-|x|)^{\gamma}(\gamma+1)^{\gamma}}(1-\frac{\gamma}{1+\gamma})$ if $t=-\ln\frac{\gamma}{(1+\gamma)(1-|x|)}$. Therefore, as $x\rightarrow\pm 1$, $V_{\gamma}^{\infty}(x)\rightarrow\infty$. On the other hand, when $\gamma=0$, for trajectories starting from $[-1,1]$, they could at least stay at their initial state, and the value is $V_{\gamma}^{\infty}(x)=|x|$, which is attained at the initial time. Therefore, the convergence is uniform, and the domain is closed. To avoid this situation, we reiterate that for R-CLVF, $\gamma>0$.
The R-CLVF value of a state $x$ captures the maximum exponentially amplified distance between the optimal trajectory (starting from this state) and the zero-level set of $\ell(x)$. If this value is finite, the optimal trajectory converges to the SRCIS under an exponential rate no slower than $\gamma$. We will build up this conclusion via the following Theorems and Lemmas.
Proposition 3
The R-CLVF is locally Lipschitz continuous for all $x\in\mathcal{D}_{\gamma}$, and the Lipschitz constant depends on the state.
Proof:
See the Appendix of [33]. ∎
Further, it is not hard to see that if the domain $\mathcal{D}_{\gamma}$ is $\mathbb{R}^{n}$, then R-CLVF is radially unbounded, i.e. $\lim_{\|x\|\rightarrow\infty}V_{\gamma}^{\infty}(x)=\infty$. This is because the R-CLVF is lower bounded by a radially unbounded function $\ell$. When the domain is an open subset of $\mathbb{R}^{n}$, we have $\lim_{x\rightarrow\partial\mathcal{D}_{\gamma}}V_{\gamma}^{\infty}(x)=\infty$.
Proposition 4
When $\mathcal{D}_{\gamma}$ is an open subset of $\mathbb{R}^{n}$, $\lim_{x\rightarrow\partial\mathcal{D}_{\gamma}}V_{\gamma}^{\infty}(x)=\infty$.
Proof:
The proof of this statement requires the R-CLVF-DPP (which has not been presented yet) and is provided in the Appendix of [33]. ∎
We now show that the R-CLVF satisfies the DPP, and is the viscosity solution to a variational inequality. These two Theorems are crucial for the proof of later results. They also serve as the theoretical foundation for computing the R-CLVF numerically. The proofs of these two Theorems are considered the main contributions of the paper and provide a means for constructing the R-CLVF for general nonlinear systems with bounded control and disturbance.
Theorem 3
(R-CLVF-DPP) For all $t\leq t+\delta\leq 0$, the following is satisfied
$V_{\gamma}^{\infty}(x)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}V^{\infty}_{\gamma}(z),\hskip 14.22636pt$
$\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u,\lambda[u]))\biggl{\}}.$ (15)
Proof:
Denote the right-hand side of equation (3) as $W(x)$. From the definition of R-CLVF, $\forall\varepsilon>0$, $x\in\mathcal{D}_{\gamma}$, $\exists t\leq 0$ s.t.
$V_{\gamma}(x,t)\leq V_{\gamma}^{\infty}(x)\leq V_{\gamma}(x,t)+\varepsilon.$ (16)
From Theorem 1, $\forall t<t+\delta\leq 0$ and $\forall x\in\mathcal{D}_{\gamma}$, define $z=\xi(t+\delta;t,x,u(\cdot),\lambda[u](\cdot))$, we have:
[…]
$e^{\gamma\delta}V_{\gamma}(\xi(t+\delta),t+\delta),\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}+\varepsilon,$
$\leq$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}V_{\gamma}^{\infty}(z),$
$\hskip 42.67912pt\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}+\varepsilon=W(x)+\varepsilon.$ (17)
On the other hand, using inequality (16) and Theorem 1, $\forall\varepsilon>0$, $\forall t<t+\delta\leq 0$ and $\forall x\in\mathcal{D}_{\gamma}$ we have:
[…]
$e^{\gamma\delta}V_{\gamma}(\xi(t+\delta),t+\delta),\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}},$
$\geq$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}(V_{\gamma}^{\infty}(z)-\varepsilon),$
$\hskip 71.13188pt\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}},$
$\geq$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}V_{\gamma}^{\infty}(z),$
$\hskip 14.22636pt\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}-e^{\gamma\delta}\varepsilon=W(x)-e^{\gamma\delta}\varepsilon.$ (18)
Combining equation (3) and (3), we show $\forall\varepsilon>0$, $\forall t<t+\delta\leq 0$ and $\forall x\in\mathcal{D}_{\gamma}$:
$W(x)-e^{\gamma\delta}\varepsilon\leq V_{\gamma}^{\infty}(x)\leq W(x)+\varepsilon,$
which completes the proof.
Theorem 4
(R-CLVF-VI viscosity solution) The R-CLVF is the solution to the following R-CLVF-VI in the viscosity sense,
$\max\biggl{\{}\ell(x)-V_{\gamma}^{\infty}(x),$ (19)
$\hskip 2.84544pt\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}V_{\gamma}^{\infty}\cdot f(x,u,d)+\gamma V_{\gamma}^{\infty}(x)\biggl{\}}=0.$
Proof:
First, define $\mathcal{F}(x,v,p):\mathcal{D}_{\gamma}\times\mathbb{R}\times\mathbb{R}^{n}\mapsto\mathbb{R}$
$\mathcal{F}(x,v,p)=\max\{\ell(x)-v,\hskip 5.0ptH(x,v,p)\},$
then the R-CLVF-VI can be written as
$\mathcal{F}(x,V_{\gamma}^{\infty}(x),D_{x}V_{\gamma}^{\infty}(x))=0.$
Follow [34],
  1. 1.
    $V_{\gamma}^{\infty}$ is a viscosity sub-solution of (19), if for any $\phi\in C^{1}(\mathcal{D}_{\gamma})$, and $x$ is a local maxima for $V_{\gamma}^{\infty}-\phi$,
    $\mathcal{F}(x,V_{\gamma}^{\infty}(x),D_{x}\phi(x))\geq 0.$ (20)
  2. 2.
    $V_{\gamma}^{\infty}$ is a viscosity super-solution of (19), if for any $\psi\in C^{1}(\mathcal{D}_{\gamma})$, and $x$ is a local minima for $V_{\gamma}^{\infty}-\psi$,
    $\mathcal{F}(x,V_{\gamma}^{\infty}(x),D_{x}\psi(x))\leq 0.$ (21)
W.O.L.G, we could always assume $x$ is a strict local maxima (minima), and the maximum (minimum) value is 00, i.e., $V_{\gamma}^{\infty}(x)-\phi(x)=0$ ($V_{\gamma}^{\infty}(x)-\psi(x)=0$). As mentioned before, the definitions of the viscosity solution for TVP and IVP are different. Here, since the R-CLVF is the limit function of TV-R-CLVF, we use the definition for the TVP.
We start with the sub-solution. Assume (20) is wrong, i.e.,
$\mathcal{F}(x,V_{\gamma}^{\infty}(x),D_{x}\phi(x))<0.$
Then there exists $\theta_{1},\theta_{2}>0$ s.t. both the followings hold
$\ell(x)-V_{\gamma}^{\infty}(x)\leq-\theta_{1}<0,$ (22)
$\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x)\cdot f(x,u,d)+\gamma V_{\gamma}^{\infty}(x)\leq-\theta_{2}<0.$ (23)
By continuity of $\ell$ and $\xi$, there exists $\delta_{1}>0$ s.t. for any $\theta_{1}>0$, $u(\cdot),\lambda$, $s\in[t,t+\delta_{1}]$
$|e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda))-\ell(x)|\leq\frac{\theta_{1}}{2}.$ (24)
Combined with (22), we have
$e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda))\leq\ell(x)+\frac{\theta_{1}}{2}\leq V_{\gamma}^{\infty}(x)-\frac{\theta_{1}}{2}.$ (25)
Further, since $V_{\gamma}^{\infty}(x)=\phi(x)$,  (23) can be written as
$\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x)\cdot f(x,u,d)+\gamma\phi(x)\leq-\theta_{2}.$
and for any $d\in\mathcal{D}$,
$\min_{u\in\mathcal{U}}D_{x}\phi(x)\cdot f(x,u,d)+\gamma\phi(x)\leq-\theta_{2}.$
Since $\phi\in C^{1}$, there exists $\delta>0,\bar{u}\in\mathcal{U}$, $s\in[t,t+\delta]$ s.t.
$D_{x}\phi(\xi(s))\cdot f(\xi(s),\bar{u}(s),\lambda[\bar{u}](s))+\gamma\phi(x(s))\leq-\frac{\theta_{2}}{2}.$
Muliplly both side with $e^{\gamma(s-t)}$ and integrate on both side for $s\in[t,t+\delta]$, we get:
$e^{\gamma\delta}\phi(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))-\phi(x)\leq-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Since $V_{\gamma}^{\infty}-\phi$ has local maxima at $x$ and the maximum value is 00, we have
$e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
$\leq$ $\phi(x)-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}=V_{\gamma}^{\infty}(x)-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$ (26)
Combine (25) and (26), and because $d$ is arbitrary, we have
$V_{\gamma}^{\infty}(x)-\min\bigl{\{}\frac{\theta_{1}}{2},\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}\bigr{\}}$
$\geq$ $\sup_{\lambda}\max\bigl{\{}e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda)),$
$\hskip 60.00009pte^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))\bigr{\}}$
$\geq$ $V_{\gamma}^{\infty}(x),$
where the last inequality is from the DPP (3). This is clearly a contradiction, so (20) must hold, i.e., $V_{\gamma}^{\infty}$ is a viscosity sub-solution.
Now, we examine the super-solution. Assume (21) is wrong, i.e.,
$\mathcal{F}(x,V_{\gamma}^{\infty}(x),D_{x}\psi(x))>0.$
Then $\exists\theta_{1},\theta_{2}>0$ s.t. one of the followings hold
$\ell(x)-V_{\gamma}^{\infty}(x)\geq\theta_{1}>0,$ (27)
$\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x)\cdot f(x,u,d)+\gamma V_{\gamma}^{\infty}(x)\geq\theta_{2}>0.$ (28)
If (27) holds, then from (24), there exists $\delta_{1}>0$ s.t. for any $\theta_{1}>0,u(\cdot),\lambda$ and $s\in[t,t+\delta_{1}]$
$e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda))\geq\ell(x)-\frac{\theta_{1}}{2}\geq V_{\gamma}^{\infty}(x)+\frac{\theta_{1}}{2}.$
Plug in the DPP (3), we have
$V_{\gamma}^{\infty}(x)\geq$ $\sup_{\lambda}\inf_{u(\cdot)}\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda))$
$\geq$ $V_{\gamma}^{\infty}(x)+\frac{\theta_{1}}{2},$
which is a contradiction. Therefore (27) cannot be true.
If (28) holds, then from the same derivation of (26), there exists $\delta_{2}>0$, $\exists\lambda,\forall u(\cdot)$, s.t. $\forall s\in[t,t+\delta_{2}]$
$e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
$\geq$ $\phi(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}=V_{\gamma}^{\infty}(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Again plug in DPP (3), we have
$V_{\gamma}^{\infty}(x)\geq$ $\sup_{\lambda}\inf_{u(\cdot)}e^{\gamma\delta}V_{\gamma}^{\infty}(\xi(\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]))$
$\geq$ $V_{\gamma}^{\infty}(x)+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma},$
which is a contradiction. Therefore (28) cannot be true. Combined, (21) must hold, so $V_{\gamma}^{\infty}$ is a viscosity super-solution.
However, it should be noted that the R-CLVF-VI (19) may have multiple solutions given different choices of $\gamma$. To see this, let’s see the 1D system $\dot{x}=u$, where $u\in[-2,2]$. It is not hard to check that for all $x>0$, the optimal control is $u^{*}=-2$ and for all $x<0$, $u^{*}=-2$. Therefore, we could easily compute the value function as
$V_{\gamma}^{\infty}(x)=\begin{cases}\frac{2}{\gamma}e^{\frac{\gamma|x|-2}{2}}&|x|>\frac{2}{\gamma}\\ |x|&|x|\leq\frac{2}{\gamma}\end{cases},$ (29)
with gradient
$\frac{d\,V_{\gamma}^{\infty}}{dx}=\begin{cases}sign(x)\cdot e^{\frac{-\gamma|x|-2}{2}}&|x|>\frac{2}{\gamma}\\ sign(x)&|x|\leq\frac{2}{\gamma}\end{cases}.$
The only non-differentiable point is $x=0$, with subdifferential $p^{-}\in[-1,1]$. It can be checked that the value function statsifies (19) in the viscosity sense. However,
$U(x)=\begin{cases}e^{\frac{\gamma(|x|-b)}{2}}&|x|>a\\ |x|&|x|\leq a\end{cases}$ (30)
is also one viscosity solution for (19), with any $a\gamma=2$, $b=\frac{2-2\ln{a}}{\gamma}$. The R-CLVF (29) is the viscosity solution for any $\gamma$.
Note that in Theorem 4, we do not specify the boundary condition. Therefore, the R-CLVF-VI cannot be directly solved. However, this is not a problem both theoretically and numerically. Theoretically, we only need to show that the R-CLVF satisfies this R-CLVF-VI in the viscosity sense. Numerically, we build up the solver based on the Level-set Toolbox [35], which is only applicable to time-varying PDEs (like equation (6) and (13)). Therefore, equation (13) is solved and backpropagated using DP until convergence, to get the R-CLVF [30].
Remark 5
The non-uniqueness of R-CLVF-VI does not affect the main result of this article. As will be shown later, the viscosity solution of R-CLVF-VI indicates the exponential stabilizability. Though the R-CLVF is not the unique one, we have proved that it is indeed one of the solutions.
Though uniqueness is not guaranteed, we can still provide the following result from Theorem 4.
Proposition 5
At any state (differentiable or non-differentiable) in the domain $\mathcal{D}_{\gamma}$ of the R-CLVF, $\forall d\in\mathcal{D}$, there exists some control $u\in\mathcal{U}$ such that
$\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}\dot{V}^{\infty}_{\gamma}\leq-\gamma V_{\gamma}^{\infty}.$ (31)
The proof can be analogously obtained from [22] and is omitted. This Proposition is vital for the proof of Lemma 6, as it states the Lie derivative of the R-CLVF along the system dynamics is less than or equal to $\gamma V_{\gamma}^{\infty}$.
III-C Existence of R-CLVF
We are now ready to present the main result of the article by introducing Lemmas 5 and 6 that will together form Theorem 7. We first show that one sufficient condition of the existence of the R-CLVF on $\mathcal{D}_{\gamma}$ is the exponential stabilizability from the ROES to the SRCIS. Later, we will show that this is also a necessary condition.
Lemma 5
The R-CLVF exists on an open set $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$) if the system is exponentially stabilizable (under rate $\gamma$) to its SRCIS from $\mathcal{D}_{\text{ROES}}$ (or $\mathbb{R}^{n}$), despite worst-case disturbance. Further $\mathcal{D}_{\gamma}=\mathcal{D}_{\text{ROES}}$.
Proof:
Assume the system is exponentially stabilizable to the SRCIS. Using the Definition 2, we have $\forall\lambda\in\Lambda$, $\exists u^{*}(\cdot)\in\mathbb{U}$, $\exists c>0$ s.t.
$dst(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]);\mathcal{I}_{m})\leq ce^{-\gamma(s-t)}dst(x;\mathcal{I}_{m}).$
Plug in equation (2),
$\min_{a\in\partial\mathcal{I}_{m}}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-a\|$
$\hskip 60.00009pt\leq ce^{-\gamma(s-t)}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|.$ (32)
Plug in $\ell(x)=h(x)-V^{\infty}_{m}=\|x\|-V^{\infty}_{m}$, we have
$\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))$
$=$ $h(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))-V^{\infty}_{m}$
$=$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\max_{a\in\partial\mathcal{I}_{m}}\|a\|$
$\leq$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\min_{a\in\partial\mathcal{I}_{m}}\|a\|,$
and
$\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\min_{a\in\partial\mathcal{I}_{m}}\|a\|$
$=$ $\min_{a\in\partial\mathcal{I}_{m}}\bigl{(}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\|-\|a\|\bigl{)}$
$\leq$ $\min_{a\in\partial\mathcal{I}_{m}}\bigl{(}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-a\|\bigl{)}$
$\leq$ $ce^{-\gamma(s-t)}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|,$ (33)
where we used equation (5) for the last inequality. Multiply $e^{\gamma(s-t)}$ on both side
$e^{\gamma(s-t)}\bigl{(}\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))\bigl{)}$
$\leq$ $e^{\gamma(s-t)}ce^{-\gamma(s-t)}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|=c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|,$
which holds for all $s\in[t,0]$. Therefore
$V_{\gamma}(x,t)=$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\left(\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))\right)$
$\leq$ $c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|.$
This upper bound $c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$ is independent of $t$, therefore as $t\rightarrow-\infty$, we have $V_{\gamma}^{\infty}(x)\leq c\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$. Since the R-CLVF monotonically increases, we conclude that the limit in (14) exists $\forall x\in\mathcal{D}_{\text{ROES}}$, and $\mathcal{D}_{\gamma}=\mathcal{D}_{\text{ROES}}$.
This Lemma shows that for nonlinear systems, the existence of the R-CLVF can be justified by analyzing the system dynamics.
One natural question to ask is, does the existence of the R-CLVF in turn imply the exponential stabilizability of the system? The answer is yes, as provided in the following Lemma.
Lemma 6
The system can be exponentially stabilized to its smallest robustly control invariant set $\mathcal{I}_{m}$ from $\mathcal{D}_{\gamma}\setminus\mathcal{I}_{m}$ (or $\mathbb{R}^{n}\setminus\mathcal{I}_{m}$), if the R-CLVF exists in $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$).
Proof:
Assume the limit in (14) exists in $\mathcal{D}_{\gamma}$. For any initial state $x\in\mathcal{D}_{\gamma}\setminus\mathcal{I}_{m}$, consider the optimal trajectory $\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])$ $\forall t\leq s\leq 0$. From Proposition 5:
$D_{x}V_{\gamma}^{\infty}(x)\cdot f(x,u^{*},d^{*})=\dot{V}_{\gamma}^{\infty}\leq-\gamma V^{\infty}_{\gamma}.$
Using the comparison principle, we have $\forall s\in[t,0]$,
$V_{\gamma}^{\infty}\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big{)}\leq e^{-\gamma(s-t)}V_{\gamma}^{\infty}(x).$ (34)
Since $V_{\gamma}(x,0)\leq V_{\gamma}^{\infty}(x)$, we have:
$\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p\|$
$\leq$ $V_{\gamma}^{\infty}\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big{)}+V^{\infty}_{m}.$
Therefore, $\forall s\in[t,0]$ we have:
$\min_{a\in\partial\mathcal{I}_{m}}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-a\|$
$=$ $\min_{a\in\partial\mathcal{I}_{m}}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p+p-a\|$
$\leq$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p\|+\min_{a\in\partial\mathcal{I}_{m}}\|p-a\|$
$\leq$ $\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-p\|+V^{\infty}_{m}$
$\leq$ $V_{\gamma}^{\infty}\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])\big{)}+2V^{\infty}_{m}.$
Plugging in (34) gives us
$\min_{a\in\partial\mathcal{I}_{m}}\|\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}])-a\|$
$\leq$ $e^{-\gamma(s-t)}V_{\gamma}^{\infty}(x)+2V^{\infty}_{m}$
$\leq$ $e^{-\gamma(s-t)}k_{1}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|+e^{-\gamma(s-t)}k_{2}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|$
$=$ $e^{-\gamma(s-t)}(k_{1}+k_{2})\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|.$
where
$k_{1}=\frac{V_{\gamma}^{\infty}(x)}{\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|},\hskip 10.00002ptk_{2}=\frac{2V^{\infty}_{m}}{e^{\gamma t}\min_{a\in\partial\mathcal{I}_{m}}\|x-a\|},$
and $0<k_{1},k_{2}<\infty$ for any given $x\notin\mathcal{I}_{m}$. In other words, the controlled system can be locally exponentially stabilized to $\mathcal{I}_{m}$ from $\mathcal{D}_{\text{ROES}}$, if the R-CLVF exists on $\mathcal{D}_{\gamma}$. Further, if the R-CLVF exists on $\mathbb{R}^{n}$, the above result holds globally.
This means for a complex nonlinear system, we could find its SRCIS, and check whether (and from where) it can be exponentially stabilized to its SRCIS by computing the R-CLVF of it. More specifically, it finds the maximum region, from where the system can be stabilized to its SRCIS under the user-specified exponential rate $\gamma$, despite worst-case disturbance.
Combining Lemma 5 and Lemma 6, we directly have the following theorem.
Theorem 7
The system can be exponentially stabilized to its SRCIS $\mathcal{I}_{m}$ from $\mathcal{D}_{\gamma}\setminus\mathcal{I}_{m}$ (or $\mathbb{R}^{n}\setminus\mathcal{I}_{m}$), if and only if the R-CLVF exists in $\mathcal{D}_{\gamma}$ (or $\mathbb{R}^{n}$).
This Theorem extends the classic ‘CLF’ results that stabilize systems to the origin in two ways. First, it is applicable to a broader class of systems (i.e., systems with no equilibrium points). Second, it guarantees the exponential rate, which is specified by the user.
III-D Impact of the Gamma
We highlight the impact of $\gamma$ in this part. The first result is that for all $\gamma\geq 0$, the zero sub-level sets of the R-CLVFs are all the same, presented in Lemma 8. This Lemma guarantees that the shape and size of SRCIS only depend on dynamics and the norm we pick, and for all different $\gamma$, we stabilize the system to the same SRCIS. Denote the zero sub-level sets of TV-R-CLVF and R-CLVF as
[…]
The TV-R-CLVF (and the R-CLVF) with different $\gamma$ has the same zero sub-level set.
Lemma 8
For all $\gamma>0$, $\mathcal{Z}_{\gamma}(t)$ are the same. Further, $\mathcal{Z}_{\gamma}^{\infty}$ are also the same and $\mathcal{Z}_{\gamma}^{\infty}=\mathcal{I}_{m}$.
Proof:
We only prove the first statement, as the second statement can be proved with the same process, and is an easy extension of the first statement. Given any $x$, assume $V_{\gamma_{1}}(x,t)<0$ and $V_{\gamma_{2}}(x,t)>0$. Since $e^{\gamma_{1}(s-t)}>0$, $\forall\lambda$, there must exist $u_{1}(\cdot)$ s.t.
$\ell(\xi(t_{1};t,x,u_{1}(\cdot),\lambda[u_{1}])<0$
for some $t_{1}$. On the other hand, since $V_{\gamma_{2}}(x,t)\geq 0$ and $e^{\gamma_{2}(s-t)}>0$, there exists $\bar{\lambda}$ and for all $u(\cdot)$ s.t.
$\ell(\xi(t_{1};t,x,u(\cdot),\bar{\lambda}[u])\geq 0$
for all $s\in[t,0]$. However, applying $u_{1}(\cdot)$ gives
$\ell(\xi(t_{2};t,x,u_{1}(\cdot),\bar{\lambda}[u_{1}])<0$
for some $t_{2}$, which is a contradiction. Therefore, $V_{\gamma_{1}}(x,t)<0$ immplies $V_{\gamma_{1}}(x,t)\leq 0$. Switching $\gamma_{1}$ and $\gamma_{2}$, we get the same result, and therefore we have $V_{\gamma_{1}}(x,t)\leq 0$ if and only if $V_{\gamma_{1}}(x,t)\leq 0$. With the same process, we could also show that $V^{\infty}_{\gamma_{1}}(x,t)\leq 0$ if and only if $V^{\infty}_{\gamma_{1}}(x,t)\leq 0$. Further, the above inequalities also hold when either $\gamma_{1}$ or $\gamma_{2}$ is 0, and when $\gamma=0$, in the infinite-time horizon, the zero sub-level set is the SRCIS. Therefore, we have proved $\mathcal{Z}_{\gamma}^{\infty}=\mathcal{I}_{m}$.
Note that when $\gamma>0$, the R-CLVF can have a negative value. Consider an initial state in the SRCIS, then $\ell(x)<0$. Then, for its value to be non-negative, $\ell(\xi(s;t,x,u(\cdot),d(\cdot)))$ has to converge to 0 with an exponential rate $\gamma$, which cannot be guaranteed. In fact, in side SRCIS, the R-CLVF is a robust control barrier-value function [29, 36] with respect to the obstacle set defined by $\ell(x)>0$.
The other question to be answered is: given that the R-CLVF with $\gamma$ exists on $\mathcal{D}_{\gamma}$, is $\gamma$ the fastest exponential rate of convergence? Revisting the example used in Remark. 4, we could see that if there exists a control signal for all possible disturbance strategies s.t. the trajectory can reach the $\mathcal{I}_{m}$ in finite time (meaning $\xi(s;t,x,u(\cdot),d(\cdot))\in\mathcal{I}_{m}$ at some $s$), the R-CLVF value at that state will be finite for all $\gamma>0$. Further, if this is the case for all states in an open set, then the R-CLVF exists for all $\gamma>0$ on that open set. This is what happened in that 1D example. On the other hand, if there exists no control signal for all disturbance strategies s.t. the trajectory reaches the $\mathcal{I}_{m}$, we cannot conclude that there exists some $\gamma_{1}>0$ s.t. $\forall\gamma>\gamma_{1}$, the R-CLVF does not exist. In general, we can only conclude the following result:
Remark 6
From (10) and (14), it can be seen that if $\gamma_{1}>\gamma_{2}$, then $V^{\infty}_{\gamma_{1}}>V^{\infty}_{\gamma_{2}}$. Assume their corresponding domain is $\mathcal{D}_{\gamma_{1}}$ and $\mathcal{D}_{\gamma_{2}}$, we have $\mathcal{D}_{\gamma_{1}}\subset\mathcal{D}_{\gamma_{2}}$. From Theorem 7, we conclude that a larger $\gamma$ corresponds to a faster convergence rate, while a smaller ROES. In other words, the user can trade off between a faster convergence rate and a larger ROES.
With all the results presented, we showed that the R-CLVF is the Lipschitz continuous viscosity solution to the R-CLVF-VI, and satisfies the dynamic programming principle, which provides the theoretical foundation for the numerical computation of the R-CLVF. We also showed that the existence of the R-CLVF is equivalent to the robust exponential stabilizability of the SRCIS of a given system. Further, we showed how the parameter $\gamma$ can affect the R-CLVF. In the next section, we provide a way to synthesize feedback controllers.
III-E R-CLVF-QP
For a control and disturbance-affine system,
$\dot{x}=f\big{(}x,u,d\big{)}=f_{x}(x)+g_{u}(x)u+g_{d}(x)d,$ (35)
where $f_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, $g_{u}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m_{u}}$, $g_{d}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m_{d}}$. Then, (31) is equivalent to a linear inequality in $u$:
$D_{x}V_{\gamma}^{\infty}(x)\cdot f_{x}(x)+\min_{u\in\mathcal{U}}D_{x}V_{\gamma}^{\infty}(x)\cdot g_{u}(x)u$
$\hskip 50.00008pt+\max_{d\in\mathcal{D}}D_{x}V_{\gamma}^{\infty}(x)\cdot g_{d}(x)d\leq-\gamma V^{\infty}_{\gamma}(x).$
Theorem 9
(Feasibility Guaranteed R-CLVF-QP) Given some reference control $u_{r}$, the optimal controller can be synthesized by the following CLVF-QP with guaranteed feasibility $\forall x\in\mathcal{D}_{\gamma}$.
$k(x)=u^{*}=\arg\min_{u\in\mathcal{U}}\quad(u-u_{r})^{T}(u-u_{r}),$
subject to $D_{x}V_{\gamma}^{\infty}(x)\cdot f_{x}(x)+D_{x}V_{\gamma}^{\infty}(x)\cdot g_{u}(x)u$
$\hskip 20.00003pt+\max_{d\in\mathcal{D}}D_{x}V_{\gamma}^{\infty}(x)\cdot g_{d}(x)d\leq-\gamma V_{\gamma}^{\infty}(x).$
Proof:
This is a direct result of Proposition 5. ∎
Note that the QP controller is only point-wise optimal, with respect to “staying close to the reference controller.” It is not optimal w.r.t. the value function. Further, since the R-CLVF is only Lipschitz continuous, its gradient may not be continuous; hence, the QP solution $u=k(x)$ is also not continuous. This may cause the solution of the closed-loop system to lose its uniqueness guarantee [37]. However, such a problem can be solved by considering the sample-and-hold solution as introduced in [38]. The sample-and-hold solution can be viewed as treating the input of the feedback law as a piecewise-continuous (in $s$) input signal, and therefore, the existence and uniqueness can be guaranteed. Further, this type of solution matches with the numerical implementation.
In fact, the relation between the stabilizability, the existence of a CLF, and the synthesis of smooth feedback controllers is quite tricky. Even if a continuously differentiable CLF is obtained, we can only guarantee to synthesize a continuous feedback controller, and the resulting closed-loop system will still face the problem of non-existence and non-uniqueness of its solution in the classic sense. Differential inclusion is another popular approach that is used to solve this issue [8].
IV NUMERICAL BENEFITS
In the numerical computation of the R-CLVF, equation (3) is solved on a discrete grid, until some convergence threshold is met, which leads to the well-known “curse of dimensionality.” In this section, we provide two main methods to overcome this issue: the warmstarting technique and the system decomposition technique. Necessary proofs are provided, and the effectiveness is validated with a 10D example in the numerical example.
IV-A R-CLVF with Warmstarting
In the previous work, we introduced a two-step process: first, the SRCIS is computed, then the R-CLVF is computed. This process requires solving the TV-R-CLVF-VI two times, each with different initializations. In this subsection, we show that the converged value function for the first step can be used to warmstart the second step computation.
Denote the TV-R-CLVF with initial value $k(x)$ as $\bar{V}_{\gamma}(x,t)$, and the infinite time value function as $\bar{V}^{\infty}_{\gamma}(x)$, with the corresponding domain $\bar{\mathcal{D}}_{\gamma}$. We only change the initial value, and still have the same loss function $\ell(x)$ for $\bar{V}_{\gamma}(x,t)$ and $V_{\gamma}(x,t)$.
Theorem 10
For all initialization $\bar{V}_{\gamma}(x,0)=k(x)$, we have $\bar{V}_{\gamma}(x,t)\geq{V_{\gamma}}(x,t)$ holds $\forall x$, $\forall t<0$.
Proof:
We show this results for three cases: $k(x)=\ell(x)$, $k(x)>\ell(x)$, and $k(x)<\ell(x)$.
(1) Assume $k(x)=\ell(x)$. Then $\bar{V}_{\gamma}(x,t)\geq{V_{\gamma}}(x,t)$ holds $\forall x$ and $t$.
(2) Assume $k(x)>\ell(x)$. From (1), $\forall t<t+\delta=0$:
$\bar{V}_{\gamma}(x,t)=$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}\bar{V}_{\gamma}(z,0),\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}$
$=$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}k(\xi(0)),\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}$
$\geq$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma\delta}\ell(\xi(0)),\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}$
$=$ $V_{\gamma}(x,t)$
(3) Assume $k(x)<\ell(x)$. Then, at time $t=0$, we have $\bar{V}_{\gamma}(x,0)<V_{\gamma}(x,0)$. Consider an infinitesimal time step $\delta t<0$ and $\delta t\rightarrow 0^{-}$, using (3), we have:
$\bar{V}_{\gamma}(x,0^{-})=$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma 0^{-}}k(\xi(0^{-})),$
$\hskip 50.00008pt\max_{s\in[0^{-},0]}e^{\gamma(s-0^{-})}\ell(\xi(s)\biggl{\}}$
$=$ $\max\biggl{\{}e^{\gamma t_{1}}k(\xi(0)),e^{-\gamma 0^{-}}\ell(\xi(0^{-})\biggl{\}}$
$=$ $e^{-\gamma 0^{-}}\ell(\xi(0^{-})$
$\geq$ $\ell(\xi(0^{-})=V_{\gamma}(x,0^{-}),$
in other words, after one infinitesimal small step, we get $\bar{V}_{\gamma}(x,t^{-})>V_{\gamma}((x,t^{-})$. Now, replace $k(x)=\bar{V}_{\gamma}(x,t^{-})$, we return to the second case, and the remaining proof follows.
Theorem 10 shows that no matter what the initial condition is, the value function propagated with this initial condition is always an over-approximation of the TV-R-CLVF. This also holds for the R-CLVF.
Proposition 6
If $\bar{V}_{\gamma}^{\infty}(x)$ exists on $\bar{\mathcal{D}_{\gamma}}$, then $\bar{V}_{\gamma}^{\infty}(x)\geq V_{\gamma}^{\infty}(x)$ and $\bar{\mathcal{D}}_{\gamma}\subseteq\mathcal{D}_{\gamma}$.
Proof:
The first part is a direct result from Theorem 10. The second part can be proved by contradiction. Assume $x\ \in\bar{\mathcal{D}}_{\gamma}$ but $x\notin{\mathcal{D}_{\gamma}}$. This means $\bar{V}_{\gamma}^{\infty}(x)$ is finite, but ${V_{\gamma}^{\infty}}(x)$ is infinite, which contradicts the first part of this proposition. ∎
The above results are inexact warmstartings, which cannot be used directly in most cases, as we want exact R-CLVF. However, they are vital for the proof of exact warmstarting, which is presented here. We show that for certain initializations, we could recover the exact R-CLVF.
Theorem 11
For initialization $\bar{V}_{\gamma}(x,0)=k(x)\leq V_{\gamma}^{\infty}(x)$, we have $\bar{V}_{\gamma}^{\infty}(x)=V_{\gamma}^{\infty}(x)$.
Proof:
Denote $\tilde{k}(x)=V_{\gamma}^{\infty}(x)$, and the value function initialized with $\tilde{k}(x)$ as $\tilde{V}_{\gamma}(x,t)$. we have $\forall x,t<t+\delta=0$:
$\tilde{V}_{\gamma}(x,t)=$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{-\gamma t}\tilde{V}_{\gamma}(z,0),\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}$
$=$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{-\gamma t}\tilde{k}(\xi(0)),\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}$
$\geq$ $\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{-\gamma t}k(\xi(0)),\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}}$
$=$ $\bar{V}_{\gamma}(x,t).$
Note that $V_{\gamma}^{\infty}(x)$ is the already the converged value function, we have $V_{\gamma}^{\infty}(x)=\tilde{V}_{\gamma}^{\infty}(x,t)\geq\bar{V}_{\gamma}(x,t)$.
Similar to Propsition 6, If $V_{\gamma}^{\infty}(x)$ exists on $\mathcal{D}_{\gamma}$, then $\bar{V}_{\gamma}^{\infty}(x)\leq{V_{\gamma}^{\infty}}(x)$, and $\mathcal{D}_{\gamma}\subseteq\bar{\mathcal{D}}_{\gamma}$. Combined, we get $\mathcal{D}_{\gamma}=\bar{\mathcal{D}}_{\gamma}$, and $\forall x\in\mathcal{D}_{\gamma}$, $\bar{V}_{\gamma}^{\infty}(x)=V_{\gamma}^{\infty}(x)$.
Algorithm 1 Obtaining the R-CLVF with warmstarting
1:: $f(x,u,d)$, $\mathcal{U}$, $\mathcal{D}$, $\gamma>0$, convergence threshold $\Delta$, $\ell(x)$, $\delta t$.
2:Output: $V_{\gamma}^{\infty}(x)$, $\mathcal{I}_{m}$
3:Initialization:
4:$V(x,t_{0})\leftarrow\ell(x)$
5:Find $\mathcal{I}_{m}$
6:$V^{\infty}(x)\leftarrow$ update_value($f$, $\mathcal{U}$, $\mathcal{D}$, $\Delta$, $\delta t$, $V(x,0)$, $\ell(x)$)
7:$V^{\infty}_{m}\leftarrow\min_{x}V^{\infty}(x)$, $\mathcal{I}_{m}\leftarrow\{V^{\infty}(x)=V^{\infty}_{m}\}$
8:Find R-CLVF
9:$\ell(x)\leftarrow\ell(x)-V^{\infty}_{m}$, $V(x,t_{0})\leftarrow V^{\infty}(x)-V^{\infty}_{m}$
10:$V_{\gamma}^{\infty}(x)\leftarrow$ update_value($f$, $\mathcal{U}$, $\mathcal{D}$, $\Delta$, $\delta t$, $V(x,0)$, $\ell(x)$)
11:update_value($f$, $\mathcal{U}$, $\mathcal{D}$, $\Delta$, $\delta t$, $V(x,0)$, $\ell(x)$)
12:$t\leftarrow 0$
13:while $dV\geq\Delta$ do
14:     $V(x,t+\delta t)\leftarrow V(x,t)$
15:     update $V(x,t+\delta t)$ using equations (1) (13)
16:     $dV=\min_{x}(V(x,t+\delta t)-V(x,t))$
17:     $t\leftarrow t+\delta t$
18:end while
Using Theorem 11, we provide an enhanced version of the original algorithm for computing the R-CLVF, shown in Alg. 1. The main difference is that after finding the SRCIS and the corresponding value function $V^{\infty}(x)$ (line 5), the next step computation (line 9) is initialized with $V^{\infty}(x)-V^{\infty}_{m}$, instead of $\ell(x)$. The exact warmstarting is guaranteed, because we can always guarantee $V^{\infty}(x)-V^{\infty}_{m}\leq V_{\gamma}^{\infty}(x)$. From the numerical examples, Alg. 1 accelerates the computation from 5% to 90%.
IV-B R-CLVF with Decomposition
To discuss the R-CLVF with decomposition, we first introduce the self-contained subsystems decomposition.
Definition 6
(Self-contained subsystem decomposition) (SCSD) Given system (1) and assume there exists state partitions $z_{1}=(x_{1},x_{c})\in\mathcal{Z}_{1}$, $z_{2}=(x_{2},x_{c})\in\mathcal{Z}_{2}$, where $x_{1}\in\mathbb{R}^{n_{1}}$, $x_{2}\in\mathbb{R}^{n_{2}}$, $x_{c}\in\mathbb{R}^{n_{c}}$, $n_{1},n_{2}>0$, $n_{c}\geq 0$, $n_{1}+n_{2}+n_{c}=n$. Assume also the control and disturbance inputs can be partitioned similarly with $v_{1}=(u_{1},u_{c})\in\mathcal{V}_{1}$, $v_{2}=(u_{2},u_{c})\in\mathcal{V}_{2}$, where $u_{1}\in\mathbb{R}^{m_{1}}$, $u_{2}\in\mathbb{R}^{m_{2}}$, $u_{c}\in\mathbb{R}^{m_{c}}$ and $m_{1}+m_{2}+m_{c}=m$. $p_{1}=(d_{1},d_{c})\in\mathcal{P}_{1}$, $p_{2}=(d_{2},d_{c})\in\mathcal{P}_{2}$, where $d_{1}\in\mathbb{R}^{p_{1}}$, $d_{2}\in\mathbb{R}^{p_{2}}$, $d_{c}\in\mathbb{R}^{p_{c}}$ and $p_{1}+p_{2}+p_{c}=p$. Given the system (1), the two subsystems of it are
$\dot{z}_{1}=f_{1}(z_{1},v_{1},p_{1}),\quad\dot{z}_{2}=f_{2}(z_{2},v_{2},p_{2}).$
Here, $x_{c}$, $u_{c}$, $d_{c}$ are called the shared state, control, and disturbances respectively.
Theorem 12
Assume the system can be decomposed into several self-contained subsystems, and there are no shared control and states between each subsystem. Denote the corresponding R-CLVFs for the subsystems as $V_{\gamma,i}^{\infty}(z_{i})$ with domain $\mathcal{D}_{\gamma_{i},z_{i}}$ and loss $\ell_{i}$, and define
$W_{\gamma}^{\infty}(x)=\max_{i}V_{\gamma,i}^{\infty}(z_{i}).$ (36)
Then, $\ell(x)=\max_{i}\ell(z_{i})$ implies $W_{\gamma}^{\infty}(x)$ is the R-CLVF of system (1).
Proof:
This is an extension of Lemma 1 of [39], and the proof can be obtained analogously. ∎
V NUMERICAL EXAMPLES
In this section, we provide three examples to showcase the main benefits of using R-CLVF: 1) it handles general nonlinear dynamics with bounded control and disturbance, 2) it finds and stabilizes the system to its SRCIS (based on different norms chosen) given a user-specified exponential rate $\gamma$, 3) with warmstarting and decomposition, the computational cost is decreased significantly. All examples are solved using MATLAB and toolboxes [40, 41]. All trajectories are generated with QP controller (9) with reference control $u_{r}=0$.
V-A 2D System Revisit
Consider again the system (8), and specify $h(x)=||x||_{\infty}$. We compute the R-CLVF with $\gamma_{1}=0.1$, $\gamma_{2}=0.3$. The results are shown in Fig. 2. It should be noted that for this system, the SRCIS for $\gamma=0.1$ and $\gamma=0.2$ are both $\mathcal{I}_{m}=\{|x|\leq 0.5,|y|\leq 0.5\}$, and ROES $\mathcal{D}_{\text{ROES}}=\{|x|>0.5,|y|<1\}\setminus\mathcal{I}_{m}$.
Figure 2: Top: R-CLVF with $\gamma=0.1$ (left) and $\gamma=0.2$ (right). Bottom left: ROES, SRCIS, and the two optimal trajectories using R-CLVF-QP controller. The ROES and SRCIS for different $\gamma$ are all the same, while the optimal trajectories are different. To see this, first consider a point on the boundary of ROES, $[0.1,1]$, $d$ will make $x$ increase 0.1 to 1, while $u$ cannot decrease $y$. Since the distance is measured by $||x||_{\infty}$, we have $\ell(\xi(s;t,x,u^{*}(\cdot),\lambda[u^{*}]))=1$, $\forall t<0$. Using equation (10) and (14), the value will be infinite. However, for any $|y|<1$, the control can decrease $y$ to 00, and for all $x$, it either goes to 0.5 or -0.5. Note both happen in a finite time horizon. Therefore, using equation (10) and (14), the value will be finite for all $\gamma\geq 0$. Bottom mid: value decay along the two optimal trajectories. All controllers were generated using R-CLVF-QP. With a 151-by-151 grid, the computation time for $\gamma=0.1$ is 215.6s with warmstarting, and 289.7s w/o warmstarting, and 211.5s with warmstarting, and 258.4s w/o warmstarting for $\gamma=0.2$.
V-B 3D Dubins Car
Consider the 3D Dubins car example:
$\dot{x}=v\cos(\theta)+d_{x},\hskip 10.00002pt\dot{y}=v\sin(\theta)+d_{y},\hskip 10.00002pt\dot{\theta}=u,$
where $v=1$ and $u\in[-\pi/2,\pi/2]$ is the control and $d_{x},d_{y}\in[-0.1,0.1]$ is the disturbance. This system has no equilibrium point. The SRCISs with different $h(x)$ are shown in Fig. 3, and the trajectory converges to the SRCIS exponentially.
Figure 3: Different SRCISs with different $h(x)$. Top left: SRCIS and optimal trajectory with $h(x)=||x||_{2}$. Top right: SRCIS and optimal trajectory with $h(x)=||x||_{Q}$, where $Q=diag[1,1,0]$. Bottom left: the value along the optimal trajectories. All controllers were generated using R-CLVF-QP. With a 51-by-51-by-53 grid, the computation time for $h(x)=||x||_{2}$ is 264s with warmstarting, and 386.6s w/o warmstarting, and 143.4s with warmstarting, and 207.7s w/o warmstarting for $h(x)=||x||_{Q}$.
Figure 4: Comparison of R-CLVF with and without warmstarting for the Z-subsystem. The difference is almost negligible.
V-C 10D Quadrotor
Consider the 10D quadrotor system:
$\dot{x}=v_{x}+d_{x},\hskip 5.69054pt\dot{v_{x}}=g\tan{\theta_{x}},\hskip 5.69054pt\dot{\theta_{x}}=-d_{1}\theta_{x}+\omega_{x},$
$\dot{\omega_{x}}=-d_{0}\theta_{x}+n_{0}u_{x},\hskip 5.69054pt\dot{y}=v_{y}+d_{y},\hskip 5.69054pt\dot{v_{y}}=g\tan{\theta_{y}},$
$\dot{\theta_{y}}=-d_{1}\theta_{y}+\omega_{y},\hskip 5.69054pt\dot{\omega_{y}}=-d_{0}\theta_{y}+n_{0}u_{y},$
$\dot{z}=v_{z}+d_{z},\hskip 5.69054pt\dot{v_{z}}=u_{z},$ (37)
where $(x,y,z)$ denote the position, $(v_{x},v_{y},v_{z})$ denote the velocity, $(\theta_{x},\theta_{y})$ denote the pitch and roll, $(\omega_{x},\omega_{y})$ denote the pitch and roll rates, and $(u_{x},u_{y},u_{z})$ are the controls. The parameters are set to be $d_{0}=10,d_{1}=8,n_{0}=10,k_{T}=0.91,g=9.81$, $|u_{x}|,|u_{y}|\leq\pi/9$, $u_{z}\in[-1,1]$, $|d_{x}|,|d_{y}|,|d_{z}|\leq 0.1$.
This 10D system can be decomposed into three subsystems: X-sys with states $[x,v_{x},\theta_{x},\omega_{x}]$, Y-sys with states $[y,v_{y},\theta_{y},\omega_{y}]$, and Z-sys with states $[z,v_{z}]$. It can be verified that all three subsystems have an equilibrium point at the origin. Further, there’s no shared control, disturbance, or state among subsystems. We use $h(x)=||x||_{\infty}$, which satisfies the condition $\ell(x)=\max_{i}\ell_{i}(z_{i})$. The R-CLVF is reconstructed using equation (36).
A comparison of the R-CLVF for the Z-sys with and without warmstarting is shown in Fig. 4, showing that the warmstarting provides the exact result. The trajectory is shown in Fig. 5.
Figure 5: Left: SRCIS of the reconstructed R-CLVF and the optimal trajectory. For Z-sys, with 101 grids for each state, time step = 0.1, convergence threshold = 0.0015, the computation time is 36.59s with warmstarting and 42.72s w/o warmstarting. For X(Y)-sys, with 17 grids for each state, time step = 0.1, convergence threshold = 0.02, the computation time is 828.27s with warmstarting and 887.79 w/o warmstarting. Right: the decay of the value along the optimal trajectory.
VI Conclusions
In this paper, we extended our preliminary work on constructing CLVFs using HJ reachability analysis to the system with bounded control and disturbances and to stabilize to a random point of interest. We provided more detailed discussions on several important Lemmas and Theorems. Additionally, warmstarting and decomposition methods are proposed to overcome the “curse of dimensionality,” and the effectiveness of both techniques is validated with numerical examples. Future directions include finding conditions under which the self-contained subsystem decomposition method provides the R-CLVF, and incorporating learning-based methods to tune the exponential rate $\gamma$ for online execution in robotics applications.
Appendix A Appendix
A-A Proof of Proposition 2
Given any open set $\mathcal{C}$ and arbitrary control and disturbance signals, as long as the solution exists, the cost function
$J_{\gamma}(t,x,u(\cdot),\lambda)$ $=\max_{s\in[t,0]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda)),$
$\leq e^{-\gamma t}\max_{s\in[t,0]}\ell(\xi(s;t,x,u(\cdot),\lambda))$ (38)
is bounded. Since this holds for arbitrary control and disturbance signals, the TV-R-CLVF is also bounded.
For the local Lipschitzness in $x$, given any $x\text{ and }y\in\mathcal{C}$, $\epsilon>0$, from (10), we have
  1. 1.
    $\exists\lambda^{*}\in\Lambda,\forall u(\cdot)\in\mathbb{U}$,
    $V_{\gamma}(x,t)\leq J_{\gamma}(t,x,u(\cdot),\lambda^{*})+\epsilon,$
  2. 2.
    $\forall\lambda\in\Lambda$, $\exists u^{*}(\cdot)\in\mathbb{U}$,
    $V_{\gamma}(y,t)\geq J_{\gamma}(t,y,u^{*}(\cdot),\lambda)-\epsilon.$
Combined, we have
$V_{\gamma}(x,t)$ $-V_{\gamma}(y,t)$
$\leq$ $\bigl{(}J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})+\epsilon\bigr{)}-\bigl{(}J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})-\epsilon\bigr{)}$
$=$ $\bigl{(}J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})-J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})\bigr{)}+2\epsilon.$
Further, we have:
$\|J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})-J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})\|$
$=$ $\|\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pt\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|$
$\leq$ $\max_{s\in[t,0]}\|e^{\gamma(s-t)}\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pte^{\gamma(s-t)}\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|$
$\leq$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\|\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pt\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|$
$=$ $e^{-\gamma t}\|\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pt\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|.$
Plug in the definition of $\ell$, we have:
$\|J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})-J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})\|$
$\leq$ $e^{-\gamma t}\bigl{\|}\|\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\|-\|\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\|\bigr{\|}$
$\leq$ $e^{-\gamma t}\|\xi(s;t,x,u^{*}(\cdot),\lambda^{*})-\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\|.$
Because of the continuous dependence on the initial condition, $\forall x,y\in\mathcal{C}$, there exists a constant $c>0$ such that
$\|\xi(s;t,x,u^{*}(\cdot),\lambda^{*})-\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\|\leq c\|x-y\|,$
Combined, we have
$\|V_{\gamma}(x,t)-V_{\gamma}(y,t)\|\leq e^{-\gamma t}c\|x-y\|+2\epsilon$ (39)
For the local Lipschitzness in $t$, for any $t<t_{1}\leq s\leq 0$, and any $\lambda,u(\cdot)$, we have
$J_{\gamma}(t_{1},x,u(\cdot),\lambda)\geq e^{\gamma(s-t_{1})}\ell(\xi(s;t_{1},x,u(\cdot),\lambda)).$
Combined with (A-A), we have
$J_{\gamma}(t,x,u(\cdot),\lambda)-J_{\gamma}(t_{1},x,u(\cdot),\lambda)$
$\leq$ $e^{-\gamma t}\max_{s\in[t,0]}\ell(\xi(s;t,x,u(\cdot),\lambda))-$
$\hskip 60.00009pte^{\gamma(s-t_{1})}\ell(\xi(s;t_{1},x,u(\cdot),\lambda))$
from (1), we have
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma(t_{1}-t)}V_{\gamma}(\xi(t_{1}),t_{1}),$
$\hskip 60.00009pt\max_{s\in[t,t_{1}]}e^{\gamma(s-t)}\ell(\xi(s))\biggl{\}},$
which means both of the following hold:
$V_{\gamma}(x,t)\geq e^{\gamma(t_{1}-t)}V_{\gamma}(\xi(t_{1}),t_{1}),$
$V_{\gamma}(x,t)\geq\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max_{s\in[t,t_{1}]}e^{\gamma(s-t)}\ell(\xi(s)).$
The first inequality and (39) implies
$V_{\gamma}(x,t)$ $\geq e^{\gamma(t_{1}-t)}V_{\gamma}(\xi(t_{1}),t_{1})$
$\geq V_{\gamma}(\xi(t_{1}),t_{1})$
$\geq V_{\gamma}(x,t_{1})-e^{-\gamma t_{1}}c\|x-\xi(t_{1})\|,$
therefore we have
$\|V_{\gamma}(x,t)-V_{\gamma}(x,t_{1})\|$ $\leq e^{-\gamma t_{1}}c\|x-\xi(t_{1})\|$
$\leq e^{-\gamma t_{1}}cL_{f}\|t_{1}-t\|.$
A-B Proof of Proposition 3
Since the convergence is point-wise, $\forall\epsilon>0$, $\exists t_{1},t_{2}<0$, s.t. $\forall x,y\in\mathcal{D}_{\gamma}$ we have
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(x)-V_{\gamma}(x,t_{1})\leq\epsilon,$
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(y)-V_{\gamma}(y,t_{2})\leq\epsilon.$
Since $V_{\gamma}(x,t)$ is non-increasing as $t\rightarrow-\infty$, taking $t_{N}=\min\{t_{1},t_{2}\}$, we have:
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(x)-V_{\gamma}(x,t_{N})\leq\epsilon,$
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(y)-V_{\gamma}(y,t_{N})\leq\epsilon.$
This gives us
$\|V_{\gamma}^{\infty}(x)-V_{\gamma}^{\infty}(y)\|\leq$ $\|V_{\gamma}(x,t_{N})-V_{\gamma}(y,t_{N})\|+2\epsilon$
$\leq$ $e^{-\gamma t_{N}}c\|x-y\|+2\epsilon.$
where we used Proposition 2 for the last inequality. Since $\epsilon$ can be chosen to be arbitrarily small, we conclude that the CLVF is locally Lipschitz in $\mathcal{D}_{\gamma}$ (refer to the proof of Theorem 3.2 of [13].) Further, given any compact subset of $\mathcal{D}_{\gamma}$, the $V_{\gamma}^{\infty}$ is Lipschitz in this compact subset.
A-C Proof of Proposition 4
Let us assume there exists a constant $C>0$ s.t. as $x\rightarrow\partial\mathcal{D}_{\gamma}$, $V_{\gamma}^{\infty}(x)\leq C$. From the R-CLVF-DPP (3), for any $t\leq t+\delta\leq 0$ both of the followings must hold:
$V_{\gamma}^{\infty}(x)\geq e^{\gamma\delta}V_{\gamma}^{\infty}(z),$
$V_{\gamma}^{\infty}(x)\geq\sup_{\lambda}\inf_{u(\cdot)}\max_{s\in[t,t+\delta]}e^{\gamma(s-t)}\ell(\xi(s;t,x,u(\cdot),\lambda)).$
From the first inequality, there exists some $\epsilon>0$ s.t.
$V_{\gamma}^{\infty}(x)\geq$ $e^{\gamma\delta}V_{\gamma}^{\infty}(z)$
$\geq$ $e^{\gamma\delta}(C-\epsilon)$
$=$ $C-(C+e^{\gamma\delta}\epsilon-e^{\gamma\delta}C).$
For any constants $C,\epsilon$, we could find a $\delta$ large enough s.t. $C+e^{\gamma\delta}\epsilon-e^{\gamma\delta}C<0$. This means
$V_{\gamma}^{\infty}(x)>C,$
which is a contradiction, and therefore such $C$ does not exists.